OpenStudy (anonymous):

(I think its -80) (fan and medal) A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, f(t), in feet, at different times, t, in seconds: f(t) = −16t2 + 48t + 100 The average rate of change of f(t) from t = 3 seconds to t = 5 seconds is _____feet per second

Mehek (mehek14):

first you have to find f(3) and f(5)

Mehek (mehek14):

tell me what (3,?) and (5,?)

OpenStudy (anonymous):

(3,2) (5,2) ? @Mehek14

Mehek (mehek14):

(5,-60) (3,100)

Mehek (mehek14):

\(\dfrac{-60-100}{5-3}=\dfrac{-160}{2}=-80\) you are correct